OpenStudy (anonymous):

How many liters of water must be added to 50L of 30% acid solution in order to produce a 20% acid solution?

OpenStudy (anonymous):

25L

OpenStudy (anonymous):

care to explain, vicky?

OpenStudy (anonymous):

yes please i dont know how you got that

OpenStudy (anonymous):

given 30% solution which means 30% of the solution is acid . so in the given 50L , about 15L is acid . we need to make it 20% , this is when 15L is the 20% of something . simple maths tells us that 20% of 75L is 15L so we need to add another 25 L to the 50L to make it 75L

OpenStudy (anonymous):

A way to approach this problem (and any similar dilution problem) is this: you are not changing the moles of acid inside the solution so you can write an equation relating the moles of acid before and after dilution:\[M_1*V_1 = M_2*V_2\]where M is the % of acid, and V is the total volume of the solution

OpenStudy (anonymous):

im so conufused

OpenStudy (anonymous):

\[30*50 = 20*(50+x)\] solving for x gives us the volume of water added in liters

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

(30/100)*50=15 (100/20)*15=75 75-50=25

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

wait so 25 is the right answer

OpenStudy (anonymous):

right

OpenStudy (anonymous):

oh ok thank you so much